Determine whether equation is independent, dependent, or inconsistent.
x - y = 3
-6 + 6y = 17

Answer by Alan3354(69443)

x - y = 3
-6 + 6y = 17 --> 6y = 23
y = 23/6
-------------
x - (23/6) = 3
x = 3 + (26/3)
x = 35/3